A company implemented an exercise program to see if it would improve job satisfaction. Job satisfaction scores for

10 randomly selected workers before and after the program are shown below. Complete parts a) through c) below.

Worker

1

2

3

4

5

6

7

8

9

10

Before

40

12

38

13

18

41

28

22

45

11

After

69

27

35

30

19

49

31

25

53

77

Compute the test statistic.

t =

Compute the P value

p=
Answer #1

From the given data A (before). alatter) 69 27 35 dA-B -29 40 -15 30. 19 Ed=-197 - Ed - 147 --14.7 in TO- S4 = Vz, Eld: - 7